A particle has such a curvilinear motion that its x coordinate is defined by π‘₯ = 5𝑑 3 βˆ’105𝑑 where π‘₯ is in inches and 𝑑 is in seconds. When 𝑑 = 2 𝑠, the total acceleration is 75 𝑖𝑛/𝑠 2 . If the y component of acceleration is constant and the particle starts from rest at the origin when 𝑑 = 0, determine its total velocity when 𝑑 = 4 𝑠.


Answer: 𝒗 = πŸπŸ–. πŸ•πŸ“ 𝒇𝒕/s


1
Expert's answer
2021-02-25T04:15:26-0500

"\begin{aligned}
&x = 5t³ - 105t\\
\\
&\dfrac{dx}{dt} = 5t² - 105\\
\\
&\therefore a_x(t) =\dfrac{d²x}{dt²} = 10t
\end{aligned}"


"a_x(2) = 10(2) = 20 \ in/s²"


"\textsf{Since, } \sqrt{(a_x)² +(a_y)²} = a"


"75 = \sqrt{20² + (a_y)²}"

"5625 = 400 + (a_y)\u00b2"

"a_y = \sqrt{5225 } = 72.28 \ in"


"v_y (t) = 72.28t\\
y(t) = 72.28t²"




"x(4) = 5(4)³ - 105(4) = 320 - 420 = -100in\\

x(0) = 5(0)³ - 105(0) = 0in"


"y(4) =72.28(4)² = 1156.5 in\\y(0) = 0in"



"r(4) = \sqrt{(x)² +(y)²} = \\\sqrt{-100² + 1156.5²} = 1160.8\ in"


"r(0) = \sqrt{(x)² +(y)²} = \\\sqrt{0² + 0²} = 0\ in"



"\textsf{Total Velocity} = \dfrac{r(4) - r(0) }{ t(4) - t(0)}
\\= \dfrac{1160.8}{4} = 290.2\ ins^{-2} = 19.8\ ft/s²"